OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 2, 6, 18, 54, ... A. an = 3(2)n-1 B. an = 2(4)n-1 C. an = 2(3)n-1

OpenStudy (anonymous):

C

OpenStudy (anonymous):



OpenStudy (anonymous):

A. \[a _{n}= 3(2)^{n-1}\] B. \[a _{n}= 2(4)^{n-1}\] C. \[a _{n}=2(3) ^{n-1}\]